OpenStudy (anonymous):

-3 + 8x - 5 = -8 solve for x

mathslover (mathslover):

\[\large{-3+8x-5=-8}\] \[\large{8x-8=-8}\] \[\large{8x-8+8=-8+8}\] \[\large{8x=0}\] \[\large{x=0}\]

mathslover (mathslover):

got it @neeedahelp ?

OpenStudy (anonymous):

yes, thanks